Exercise 12-9 During 2013, Pina Corporation spent $159,840 in research and development costs. As a result, a new product called the New Age Piano was patented. The patent was obtained on October 1, 2013, and had a legal life of 20 years and a useful life of 10 years. Legal costs of $32,400 related to the patent were incurred as of October 1, 2013 Your answer is correct. Prepare all journal entries required in 2013 and 2014 as a result of the transactions above. (Round answers to O decimal places, e.g. 8,564. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry" for the account titles and enter 0 for the amounts.) Date Account Titles and Explanation 2013Research and Development Expense Debit Credit 159840 Cash 159840 To record research and development expenses) 2013 Patents 32400 Cash 32400 (To record legal expenses) 2013 Amortization Expense 810 Patents 810 (To record amortization expense)
2014 Amortization Expense 3240 Patent 3240 SHOW LIST OF ACCOUNTS SHOW SOLUTION SHOW ANSWER LINK TO TEXT LINK TO TEXT Your answer is partially correct. Try again. On June 1, 2015, Pina spent $24,480 to successfully prosecute a patent infringement suit. As a result, the estimate of useful life was extended to 12 years from June 1, 2015. Prepare all journal entries required in 2015 and 2016. (Round answers to 0 decimal places,e.g. 8,564. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.) Credit Date Account Titles and Explanation Debit 2015 Patents 24480 Cash 24480 (To record patents) 2015 Amortization Expense 3984
To record amortization expense) 2016 Amortization Expense 4515 Patents 4515 SHOW LIST OF ACCOUNTS LINK TO TEXT LINK TO TEXT Your answer is partially correct. Try again. In 2017, Pina determined that a competitor's product would make the New Age Piano obsolete and the patent worthless by December 31, 2018. Prepare all journal entries required in 2017 and 2018. (Round answers to 0 decimal places, e g 8,564 Credit account titles are automatically inde ted when amount is entered. Do not indent manua、rno entry s required, select No Entry" for the account titles and enter 0 for the amounts.) Debit Account Titles and Explanation Credit Amortization Expense 25710 Patents 25710
Answer #1

Solution:

Question b:

Journal entries

Date Particulars Debit($) Credit($)
2015 Patents A/c Dr $24,480
To Cash A/c $24,480
2015 Amortization expense A/c Dr $3,853
To Patents A/c (1,350 + 2,503) $3,853
2016 Amortization expense A/c Dr $4,290
To Patents A/c (51,480 / 12) $4,290

Working notes:

Jan 1 to June 1:

= (32,400 / 10) * 5/12

= 1,350

June 1 to Dec 31:

= (32,400 - 810 - 3,240 - 1,350) + 24,480

= 51,480

= (51,480 / 12) * 7 / 12

= 2,503

Question c:

Journal entry

Date Particulars Debit($) Credit($)
2017 & 2018 Amortization expense A/c Dr $21,669
To Patents A/c (43,337 / 2) $21,669

Working notes:

= 51,480 - 3,853 - 4,290

= 43,337

Note:

If you are satisfy with this solution please give a thumbs up. Feel free to ask you have any doubt.

Thank you.